Laurence is painting the walls and ceiling of a rectangular bedroom that is 3 m by 4.5 m. The walls are 2.5 m high.
There is a window measuring 1 m by 0.8 m on one wall and a door that is 1 m by 2.3 m in another wall.
If a can of paint covers 20 m2 and costs $38.95, what will it cost Laurence for paint for this bedroom
1
Expert's answer
2011-12-09T09:14:53-0500
S = 3*4.5*2.5 - 3*4.5 - 1*0.8 - 1*2.3 = 3*4.5*1.5 - 0.8 - 2.3 = 17.15 ft2
So
he needs 1 can of paint to buy and will pay $38.95